OpenStudy (anonymous):

find an equation of the line containing the given pair of points. (4,2) (12,7) y=

geerky42 (geerky42):

y = mx+b Where m is slope and b is y-intercept. You can obtain m by using this:\[m = \dfrac{y_2-y_1}{x_2-x_1}\] from points \((x_2,y_2)\) and \((x_1,y_1)\) Then for b, plug in any point then solve for b.

OpenStudy (anonymous):

y-2=5/8(x-4)

OpenStudy (anonymous):

how is this y=mx+b form?